The slope and y-intercept for 4x-2y=10

Answer by benni1013(206)

Step 1: Move 4x to the other side.
-2y=10-4x
Step 2: Divide each side by -1
2y=-10+4x
Step 3: Divide each side by 2
y=-5+2x
Step 4: Rearrange
y=2x-5
The above is now in slope-intercept form. y=mx+b. m is the slope and b is the y intercept. In this case the slope is 2 and the y intercept is -5.